### v1

@@ 3.2.2. $\alpha_{u}=\alpha_{v},\beta_{u}=\beta_{v}$
We now index blocks of 0’s in $u_{long}^{+\alpha,\beta}$ clockwise and split all blocks into groups similarly to the case 3.2.2.: each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. If the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in different groups, then we can proceed with the proof as in case 3.2.2.. Assume that blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the same group, and the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ have indices $i$ and $i+hr_{2}$, respectively. If $h\neq t-1$ then we can proceed with the proof as in case 3.2.3.. Assume $h=t-1$.
From each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from every second block (from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$). If $t$ is odd, then from the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add to $u_{1}$ block $0^{\alpha_{u}}$ and blocks $0^{x}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(t-2)r_{2}$. If $t$ is even, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add to $u_{1}$ block $0^{\alpha_{u}}$ and blocks $0^{x}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(t-3)r_{2},i+(t-2)r_{2}$. If $t\neq 4$ then it is not difficult to verify that we add at most $\frac{2}{3}t$ blocks to $u_{1}$ from each group. If $t=4$, then there are at least two groups of blocks since $a\geqslant 3$ (there are at least three big blocks in $u$). So from each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we add $\frac{t}{2}=2$ blocks to $u_{1}$ and from the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we add $3$ blocks. So we add at most $\frac{2}{3}(a+2)$ blocks $0^{x}$ to $u_{1}$. We let $u_{1}^{\prime}$ denote the obtained subword of $u$. The proof in this case can be completed similarly to the case 3.2.2.
The lemma is proved. ∎

### v2

@@ 4.2 Lemmas constituting the proof
We now index blocks of 0’s in $u_{\operatorname{long}}^{+\alpha,\beta}$ clockwise and split all blocks into groups similarly to Case 3.2.2: each group contains $t$ blocks with indices $i,i+r_{2},i+2r_{2},\ldots,i+(t-1)r_{2}$ for some $i$. If the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in different groups, then we can proceed with the proof as in Case 3.2.2. Assume that blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ are in the same group, and the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ have indices $i$ and $i+hr_{2}$, respectively. If $h\neq t-1$ then we can proceed with the proof as in Case 3.2.3. Assume h=t1h=t-1.
From each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add $0^{x}$ to $u_{1}$ from every second block (from blocks with indices $i,i+2r_{2},i+4r_{2},\ldots$). If $t$ is odd, then from the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add to $u_{1}$ block $0^{\alpha_{u}}$ and blocks $0^{x}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(t-2)r_{2}$. If $t$ is even, then for the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we can add to $u_{1}$ the block $0^{\alpha_{u}}$ and the blocks $0^{x}$ from blocks with indices $i+r_{2},i+3r_{2},\ldots,i+(t-3)r_{2},i+(t-2)r_{2}$. If $t\neq 4$ then it is not difficult to verify that we add at most $\frac{2}{3}t$ blocks to $u_{1}$ from each group. If $t=4$, then there are at least two groups of blocks since $a\geqslant 3$ (there are at least three big blocks in $u$). So from each group which does not contain the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we add $\frac{t}{2}=2$ blocks to $u_{1}$ and from the group containing the blocks $0^{\alpha_{u}}$ and $0^{\beta_{u}}$ we add $3$ blocks. So we add at most $\frac{2}{3}(a+2)$ blocks $0^{x}$ to $u_{1}$. We let $u_{1}^{\prime}$ denote the subword obtained from $u$. The proof in this case can be completed similarly to Case 3.2.2. ∎